The speed of light in vacuum is 3*100000000. How many nano second does it taketo travel one meter in vacuum

Grade:11

1 Answers

Arun
6 years ago
In 1 sec light travels 3 *10^8 metres
That means 
3* 10^8 m = 1 sec
Hence 1 m = 1/3*10^8 sec
 = 0.33 *10^-8
= 3.33 * 10^-9 sec
= 3.33 nanosec